A very small helium-filled balloon has a volume of 530.0 mL. At the point you released the balloon, the pressure was 756.8 mmHg at 21.69ºC. What is the volume of the balloon (in liters) once it has ascended into the stratosphere where the pressure is 0.890 mmHg and a temperature of -71.37ºC?
1
Expert's answer
2018-12-07T03:55:34-0500

Answer on Qusetion #83635, chemistry/General Chemistry

A very small helium-filled balloon has a volume of 530.0 mL. At the point you released the balloon, the pressure was 756.8 mmHg at 21.69ºC. What is the volume of the balloon (in liters) once it has ascended into the stratosphere where the pressure is 0.890 mmHg and a temperature of -71.37ºC?

Solution

Use Combined and Ideal Gas Laws to answer this question:

P1V1/T1 = P2V2/T2

where

P1 = 756.8 mmHg

V1 = 0.5300 L

T1(K) = 21.69C + 273.15 = 294.84 K

P2 = 0.890 mmHg

T2(K) = -71.37C + 273.15 = 201.78 K

756.8 mmHg *0.5300 L/294.84 K = 0.890 mmHg *V2 /201.78 K

V2 = 308.4 L

Answer: 308.4 L